OpenStudy (anonymous):

Which expressions are equal to 6 when x = 8 and y = 1? Choose all answers that are correct. A. x – y + 1 B. (x ÷ y) – 2 C. x + y – 5 D. xy – 2

OpenStudy (anonymous):

okay, we know that A is incorrect since 8−1+1=8

OpenStudy (anonymous):

we know B is correct since 8 divided by 1-2=6

OpenStudy (anonymous):

we know C is incorrect since 8+1−5=4

OpenStudy (anonymous):

we know D is correct since 8*1−2=6

OpenStudy (anonymous):

so B and D are correct